the profit obtained by manufacturing and selling x item is given by P (x) = 60x - x ^ 2. Determine the number of units that must be produced and sold in order to maximize profit.
Answer #1

Given profit function P(x) = 60x - x^2

So to get units of maximum profit, differentiating this function with respect to x

dP(x)/dx = 60 - 2x = 0

So, x = 30 units

So, profit = 60*30 - 30^2 = 900

the number of units that must be produced and sold in order to maximize profit is 30 units